# Climate in a Bottle: Towards a Generative Foundation Model for the Kilometer-Scale Global Atmosphere

### v1

@@ Appendix B Diurnal Cycle: Details
Figure SI1: Diurnal cycle diagnostics for surface air temperature. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
Appendix C Architecture: Details
Figure SI2: Position and calendar embedding.
Figure SI3: Schematic showing the UNet architecture for spherical geometry with HEALPix grids. Each convolution is applied on HEALPix faces padded as in [Karlbauer et al., 2024].
Figure SI4: Training workflow of the multi-diffusion framework. Global maps are decomposed into patches for parallel training of local denoisers conditioned on spatially aligned context.
Figure SI5: Illustration of the aggregated local denoiser replacing the global denoiser in the sampling algorithm.
Appendix D Datasets: Details
Appendix E Pseudo-code for masking in the loss function
D_yn = net(where(mask, y + n, 0), sigma)
loss = weight * ((D_yn - where(mask, y, 0)) ** 2)
loss = where(mask, loss, 0) / mask.mean()
Appendix F Super-resolution results: Details
Figure SI6: Spherical power spectra and distribution plots of ICON HPX64 super-resolution inferences across 32 validation samples. (a), (c): power spectra for 2-meter temperature and precipitation. (b), (d): log-scale distributions of 2-meter temperature and precipitation.
Super-resolution probability distribution
@@ Appendix F Super-resolution results: Details
Figure SI6 shows the probability distributions of surface air temperature and precipitation aggregating over all temporal and spatial locations. The super-resolution (red line) captures the shape of temperature and precipitation PDFs (Figure SI6a,c) including the correct power law regulating the heavy tail of low-likelihood extreme precipitation occurrence (Figure SI6c).).
Cascaded generations probability distribution
Figure SI7 shows the probability distributions computed using 64 randomly selected samples spanning from March 6, 2024, to July 23, 2025. Overall, the distributions match well, though the cascaded model struggles to produce the coldest temperatures (less than -60 deg C). The super-resolution did not have this problem, which suggests that the bias likely comes from the coarse-resolution generations in this cascaded setup.
@@
Figure SI7: Spherical power spectra and distribution plots of cascaded cBottle inferences across 64 validation samples. (a), (c): power spectra for 2-meter temperature and precipitation. (b), (d): log-scale distributions of 2-meter temperature and precipitation. Samples from cascaded and ICON are not matched.
Multidiffusion Parameter Ablation
@@
To investigate this trade-off, we evaluate various overlap sizes using the median absolute horizontal Sobel gradient of surface temperature, as shown in Figure SI8. This metric serves as a proxy for boundary artifacts, where higher gradient magnitudes indicate sharper discontinuities. To generate the plot, we aggregate all faces from 16 randomly sampled outputs. A Sobel-X filter is applied to the surface temperature field, and the median of the resulting gradient magnitudes is computed across the face, sample, and vertical dimensions.
@@
Figure SI8: Severity of patch artifacts across different overlapping pixel sizes, visualized using the mean absolute Sobel-X gradient of surface temperature field as a function of horizontal pixel position.
Figure SI9: Emulation of native-grid ICON with ERA5: Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples. All datasets are regridded to a Lambert-Conformal projection for comparison. ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps. ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
Appendix G Further climate diagnostics
Figure SI10: Seasonal cycle for (JJA- DJF) for ERA5 and cBottle. Panels are shown for Z500, surface air temperature, precipitation, and sea ice concentration.
G.1 Zonal Mean Precipitation
Figure SI11 examines mean precipitation in cBottle ERA5 and its relationship with latitude computed. Overall cBottle ERA5 captures the existence and correct positioning of local maxima associated with midlatitude storm tracks and intertropical convergence zones straddling the equator. Local imperfections include positive biases in the northern (+4%) and southern (+11%) tropical rain bands, and as much as 11% underestimation of midlatitude rainfall. These differences can in part be explained partially by sampling error.
Figure SI11: Zonal Average of precipitation ICON, ERA5, and their emulators. 95% confidence intervals for the validation set means are estimated by bootstrapping over monthly averages. Only data from the validation periods for ICON and ERA5 are used.
G.2 Power-spectral fidelity of the coarse-resolution generator
Figure SI12: (left) Spherical power spectra of versus the truth. Averages over 100 random samples from the test period are shown. The dotted lines indicate the range of noise levels used for the sampling $\sigma_{\{min,max\}}^{2}\frac{4\pi}{n_{pix}}$. The factor of $\frac{4\pi}{n_{pix}}$ scales the pixel-wise variance to the same units as the power spectra, which is variance per surface area of the unit sphere.
Figure SI13: PDF of instantaneous precipitation. Computed for same samples as in SI12.
Figure SI12 shows that the spectra of the generated coarse samples are consistent with their ICON and ERA5 ground truth. For some fields, the ICON data shows significantly more power than ERA5 at large wave numbers, but we have not confirmed if this comes from the aliasing of smaller scale modes due to the block averaging used to coarsen the native HPX1024 outputs for ICON or an actual difference.
G.3 Submonthly variance
Figure SI14: Sub-monthly standard-deviation (variance from the monthly mean averaged over all months in the test period), surface air temperature, and precipitation. Panels are shown for the ground truth and cBottle for ICON and ERA5.
Figure SI14 maps summarize the pointwise variance of samples after removing the seasonal cycle, i.e. anomalized with respect to their monthly climatology (Figure SI14). This summarizes dispersion produced by a variety of sources including diurnal and synoptic variability. Variance of 500 hPa geopotential height (top) is maximized over the Southern Ocean, North Atlantic Sea and northeast Pacific due to midlatitude weather systems, but avoids the tropics due to the weak temperature gradient constraint there [Sobel et al., 2001]. Midlatitude rainfall variance occurs over the storm tracks including the Gulf Stream and Kuroshio currents. Rainfall variance is especially high within the intertropical convergence zones. In the special case of surface temperature, diurnal variability leads to a distinct land-sea contrast – a result of solar heating causing strong day-night temperature cycles over land, where surface specific heat capacity is relatively low.
The main point of Figure SI14 is that cBottle-ERA5 and cBottle-ICON faithfully generate all of these pointwise variance features summarizing diurnal-to-synoptic variability – and with sufficient precision to detect differences between them, such as the “double ITCZ” – which manifests in cBottle-ICON and cBottle-ICON-Target as a zonal band of excessive dispersion in tropical rainfall in the eastern Pacific just south of the equator.
Appendix H Supplementary Figures
Figure SI15: Bias correction of the total cloud liquid water. (top) original ICON data and (bottom) bias corrected.

### v2

@@ B.1 Diurnal Cycle: Details
a. ERA5 b. cBottle
Refer to caption Refer to caption
c. ERA5 d. cBottle
Refer to caption Refer to caption
@@
Figure SI4: Diurnal cycle diagnostics for surface air temperature. (top row) Amplitude of the diurnal cycle. (bottom row) signal to noise ratio.
a. ERA5 b. cBottle
Refer to caption Refer to caption
c. ERA5 d. cBottle
Refer to caption Refer to caption
Figure SI5: Same as Figure SI4 but for total surface precipitation.
B.2 ENSO Response
Figure SI6: El Niño-Southern Oscillation. Regression coefficient of the Nino3.4 index against monthly average precipitation. (left) ERA5. (right) Climate in a bottle. Computed over 1990-2020 inclusive.
Figure SI6 demonstrates that cBottle produces realistic ENSO precipitation variability, measured as the regression of its generated precipitation against the Nino3.4 index associated with its input SST. Within the tropics, increases of Central Pacific precipitation during the positive phase of ENSO associate with reductions of generated precipitation over tropical Africa, Indonesia and South America, consistent with Walker Cell modulation. These results add to the evidence of cBottle’s ability to condition generated variability appropriately on SST, in addition to time of year and time of day.
B.3 Interactivity: Details
Tropical Cyclones can in principle be requested in any location, and arbitrarily many can be requested in a single sample. Some such guidance inputs may be unphysical, such as TCs on the equator or at high latitudes. We find that while the guided model can produce such unphysical samples, TCs there tend to be more muted than physical ones, as demonstrated in Fig. SI7a). Here, we guide the same way as before, but with two locations on the equator and one off the coast of Greenland. All three locations show visible pressure minima, albeit somewhat weaker in the equator cases than in Fig. 3f).
The guidance could also lead to a loss of variability in the output states. In the limit of very strong guidance and many locations, we cannot expect large variability in samples. However, in the moderate guidance setting described here, variability qualitatively seems to be preserved: In Fig SI7b) the ensemble standard deviations of a guided and unguided ensemble are similar (if the variability of the guided samples was very low, we would see a mostly blue map).
a. Unphysical samples b. Variability
Refer to caption Refer to caption
Figure SI7: Realism of generated TC states. a) The same Figure as Fig. 3f), but with TC guidance applied in different location. b) Difference between guided (TC guidance applied in one location near Florida) and unguided standard deviation in Z500 of a 10 member ensemble of states.
B.4 ICON in a bottle
B.4.1 Super-Resolution
Figure SI2 provides a first qualitative assessment of the super-resolution on its own (starting from low-resolution ICON data). Here we choose a convenient time sample from the test set, in which a West Pacific tropical cyclone occurred adjacent to a large midlatitude weather system to its northeast. The performance of super-resolution is shown for two field variables – upwelling solar radiation (first row) and precipitation rate (second row). Compared to the ground truth (right column), the results (middle column) are generally encouraging. For the tropical cyclone, appropriate orographic maximization of convection over the interior of Taiwan – not present in the macroscale inputs (left column) – is realistically generated. So too are details of the eyewall structure and surrounding rain bands, albeit with less spatial coherence than in the ground truth. Meanwhile, for the midlatitude weather system, a plausible fine-scale frontal band is generated, extending from the Kiril islands to just east of the Kamchatka peninsula, with consistent horizontal scale and position as the ground truth, across both field variables.
A more quantitative assessment is shown in Figure SI8 from the view of distributions and spectra. The plots are generated using 32 randomly selected samples spanning from March 6, 2024, to July 23, 2025. The spherical power spectra are computed by averaging the squared magnitudes of the spherical harmonic coefficients over the azimuthal index $m$ and across the temporal dimension. Correct distribution of variance as a function of spatial scale is confirmed in the power spectra for temperature and precipitation.
Interestingly, similar quality results can be obtained when only training on small 4-week subset of the 5 year ICON dataset and patching artifacts can be mitigated by the choice of the patching overlap (see Appendix B.7).
@@ B.4.1 Super-Resolution
Figure SI8(a,c) shows the probability distributions of surface air temperature and precipitation aggregating over all temporal and spatial locations. The super-resolution (red line) captures the shape of temperature and precipitation PDFs including the correct power law regulating the heavy tail of low-likelihood extreme precipitation occurrence (Figure SI8c).).
@@
Figure SI8: Spherical power spectra and distribution plots of ICON HPX64 super-resolution inferences across 32 validation samples. (a), (c): power spectra for 2-meter temperature and precipitation. (b), (d): log-scale distributions of 2-meter temperature and precipitation.
B.4.2 End-to-end synthesis: Coarse generation + Super-resolution
We now turn to results from our capstone task of attempting to generate 12M-pixel multivariate ICON states purely from noise, time of year, time of day, and a monthly mean map of SST. The left column of Figure SI9 provides a qualitative demonstration of the results (see also Figure 2). Results are shown for an arbitrary subregion spanning the entire North Pacific ocean, across four arbitrary time periods having distinct weather states. It is important to note that the ground truth (corresponding right columns) should not be compared to the model output for any matched spatial positioning of individual weather events, since the task is unconditional generation. That is, the ground truth also has distinct weather states. Rather, what is noteworthy in the comparison is the common modulation of similar microscale weather detail by variably positioned macroscale weather events.
Figure SI9: Demonstration of 2 HPX1024 samples from cascaded cBottle, showing outgoing shortwave radiation, and precipitation flux. Samples from cascaded and ICON are not matched.
Figure SI10 presents the probability distributions and spherical power spectra computed from 64 randomly selected samples generated by the cascaded cBottle model, spanning the period from January 3, 2020, to July 23, 2025. Spectra analysis again quantitatively confirm the finding. On the one hand, secondary imperfections are introduced that were not apparent when super-resolution is used in isolation of macro-scale generation (Figures SI2) such as an increased presence of high-frequency components in the precipitation spectrum. On the other hand such imperfections are minor such that the results are overall of reasonable enough quality so as to suggest the viability of cascaded super-resolution as a means to generate multivariate ICON states.
In general, the model-generated distributions exhibit good agreement with the target, though the cascaded approach shows a tendency to exaggerate extreme temperature values and slightly overpredict precipitation. The super-resolution (Figures SI2) did not have these problems, which suggests that the bias likely comes from the coarse-resolution generations in this cascaded setup.
Figure SI10: Distribution and spherical power spectra plots of cascaded cBottle inferences across 64 validation samples. Column 1: log-scale distributions of 2-meter temperature and precipitation. Column 2: power spectra for 2-meter temperature and precipitation.
B.5 Multimodal: ICON-Informed Downscaling of ERA5
Although the super-resolution model is trained exclusively on ICON data, it demonstrates zero-shot generalization capability when applied to data from other sources. In this subsection, we assess its skills on ERA5 data regridded to HPX64 resolution. The corrupted radiation channels in ERA5 have been infilled using the method described in Section 3.4. Please note that this method differs slightly from the translate-then-super-resolve approach in Figure 5h, which translates ERA5 to ICON and then super-resolves instead. The resulting super-resolved outputs are compared against both the 3-km High-Resolution Rapid Refresh (HRRR) dataset over CONUS and the native-resolution ICON model at HPX1024. All datasets used in this evaluation are regridded to a common 3-km Lambert conformal projection over the CONUS region to ensure fair spatial alignment.
Figure SI11: Illustration of the downscaled ERA5 output at 15:00 UTC on December 11, 2018, with radiative channels infilled using the cBottle HPX64 generator. The first three columns correspond to the same timestamp, while the rightmost column (ICON) shows the free-run forecast for 15:00 UTC on December 11, 2024. From top to bottom, the rows represent 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation.
Figures SI11 presents a qualitative comparison of three representative variables: 10-meter eastward wind velocity, outgoing shortwave radiation, and precipitation rate. The figure shows spatial subsets spanning the Midwest to the Eastern United States, with four columns corresponding to the original ERA5 HPX64 input, the super-resolved ERA5 output generated by our model, the HRRR reference, and the ICON HPX1024 reference. Note that the ICON simulation is a free-running model and is not temporally aligned with the ERA5 or HRRR data; however, it still serves as a useful reference for assessing small-scale structural similarity. While the large-scale patterns in the super-resolved ERA5 resemble those in the original ERA5 and HRRR data, the finer-scale structures more closely resemble those in the native-resolution ICON fields. This suggests that the model effectively emulates high-frequency features in ICON and is capable of generating high-resolution patterns and dynamics that are not explicitly present in the coarse-resolution data.
Figure SI12: Emulation of native-grid ICON with ERA5: Power spectra of cBottle-resolved ERA5 and various baselines, computed over 10 samples. All datasets are regridded to a Lambert-Conformal projection for comparison. ERA5, cBottle-resolved ERA5, and HRRR share an identical set of evaluation timestamps. ICON HPX64 and ICON HPX1024 use matching times of day and days of the year, but from different years.
The zonal-mean spherical power spectra for four representative variables are shown in Figure SI12: 2-meter temperature, outgoing shortwave radiation, and precipitation. Comparisons are made across five datasets: the ERA5 HPX64 input, the super-resolved ERA5 output generated by our model (cBottle), the HRRR reference, ICON at HPX64, and ICON at HPX1024 resolution.
B.6 Further climate diagnostics
Figure SI13: Loading pattern for the Northern Annular Mode computed from ERA5. This is the first empirical orthogonal function of the ERA5 Z1000 monthly anomalies computed over 1970-2010, north of 20 degrees N, and for the winter months of November-April.
Figure SI14: Cycle of sea ice concentration for ERA5 emulation and ground truth, computed over the test period
Figure SI15: Seasonal cycle for (JJA- DJF) for ERA5 and cBottle. Panels are shown for Z500, surface air temperature, precipitation, and sea ice concentration.
B.6.1 Zonal Mean Precipitation
Figure SI16 examines mean precipitation in cBottle ERA5 and its relationship with latitude computed. Overall cBottle ERA5 captures the existence and correct positioning of local maxima associated with midlatitude storm tracks and intertropical convergence zones straddling the equator. Local imperfections include positive biases in the northern (+4%) and southern (+11%) tropical rain bands, and as much as 11% underestimation of midlatitude rainfall. These differences can in part be explained partially by sampling error.
Figure SI16: Zonal Average of precipitation ICON, ERA5, and their emulators. 95% confidence intervals for the validation set means are estimated by bootstrapping over monthly averages. Only data from the validation periods for ICON and ERA5 are used.
B.6.2 Climate change
a
b
c
Figure SI17: Climate change. Time series of global mean anomalies from ERA5’s monthly climatology computed over 1980-2010. (a) surface air temperature, (b) precipitation, (c) 500-mb geopotential. 95% confidence-intervals for the annual mean anomaly are computed by bootstrapping over the monthly averages.
B.6.3 Power-spectral fidelity of the coarse-resolution generator
Figure SI18: (left) Spherical power spectra of versus the truth. Averages over 100 random samples from the test period are shown. The dotted lines indicate the range of noise levels used for the sampling $\sigma_{\{min,max\}}^{2}\frac{4\pi}{n_{pix}}$. The factor of $\frac{4\pi}{n_{pix}}$ scales the pixel-wise variance to the same units as the power spectra, which is variance per surface area of the unit sphere.
Figure SI19: PDF of instantaneous precipitation. Computed for same samples as in SI18.
Figure SI18 shows that the spectra of the generated coarse samples are consistent with their ICON and ERA5 ground truth. For some fields, the ICON data shows significantly more power than ERA5 at large wave numbers, but we have not confirmed if this comes from the aliasing of smaller scale modes due to the block averaging used to coarsen the native HPX1024 outputs for ICON or an actual difference.
B.6.4 Variance maps
Figure SI20: Total standard deviation maps with respect to the annual mean computed over the train and test periods for surface air temperature, and precipitation. Panels are shown for the ground truth and cBottle for ICON and ERA5. The globally-averaged RMS is shown in the label for each subpanel.
Figure SI20 maps summarize the pointwise variance of samples after removing the time mean (Figure SI20). This summarizes dispersion produced by a variety of sources including diurnal, synoptic variability, seasonal, and inter-seasonal. Variance of 500 hPa geopotential height (top) is maximized over the Southern Ocean, North Atlantic Sea and northeast Pacific due to midlatitude weather systems, but avoids the tropics due to the weak temperature gradient constraint there [53]. Midlatitude rainfall variance occurs over the storm tracks including the Gulf Stream and Kuroshio currents. Rainfall variance is especially high within the intertropical convergence zones. In the special case of surface temperature, diurnal variability leads to a distinct land-sea contrast – a result of solar heating causing strong day-night temperature cycles over land, where surface specific heat capacity is relatively low.
The main point of Figure SI20 is that cBottle-ERA5 and cBottle-ICON faithfully generate all of these point-wise variance features summarizing all sources variability – and with sufficient precision to detect differences between them, such as the “double ITCZ” – which manifests in cBottle-ICON and cBottle-ICON-Target as a zonal band of excessive dispersion in tropical rainfall in the eastern Pacific just south of the equator.
B.7 Ablations
Training data size
Figure SI21 illustrates the spherical power spectra of inference results from two models trained on datasets of differing sizes: one with 3.75 years of distinct samples (72,360 in total), and another with 4 weeks of distinct samples (1,344 in total), with each week randomly selected from a different season to ensure representative temporal coverage. Both models were trained using 95,000 total samples via repetition to match training iterations.
Figure SI21: Training data size ablation study: Power spectra from global super-resolution models trained on 3.75 years versus 4 weeks of timesteps.
Noise schedule
Figure SI22: Four consecutive frames of reflective visible light (rsut) from a sequence generated by video diffusion models trained with either a log-uniform (top) or log-normal (bottom) noise schedule. Generations are conditioned on the calendar embedding and monthly average SST. Red × marks indicate the subsolar point at each timestep.
Figure SI22 demonstrates the impact of noise schedule choice on diffusion models for atmospheric data. The bottom row shows results using a log-normal schedule ($P_{mean}=-1.2$, $P_{std}=1.2$) typically used for image generation [9]. This schedule introduces insufficient noise to mask the dominant diurnal cycle of outgoing shortwave radiation during training. Consequently, the model learns to detect and amplify residual signal patterns from noised inputs rather than using the time-of-day conditioning. However, during inference where starting from pure noise provides no such signal to extract, the model produces physically inconsistent results. Panel (g) shows this most clearly, with a “double sun” artifact with radiation appearing in multiple incorrect locations.
In contrast, the log-uniform schedule (top row) with higher $\sigma_{max}$ provides sufficient noise to completely obscure the diurnal pattern during training, forcing the model to learn the relationship between time-of-day conditioning and the spatial distribution of solar radiation. This yields physically consistent results qualitatively with correctly positioned subsolar points. This visual comparison confirms the earlier theoretical argument (see SubsectionA.3.1) that atmospheric variables, some with planetary scale patterns, require substantially higher noise ceilings than is standard for diffusion to avoid signal leak bias.
Multidiffusion overlap
@@ B.7 Ablations
To investigate this trade-off, we evaluate various overlap sizes using the median absolute horizontal Sobel gradient of surface temperature, as shown in Figure SI23. This metric serves as a proxy for boundary artifacts, where higher gradient magnitudes indicate sharper discontinuities. To generate the plot, we aggregate all faces from 16 randomly sampled outputs. A Sobel-X filter is applied to the surface temperature field, and the median of the resulting gradient magnitudes is computed across the face, sample, and vertical dimensions.
@@
Figure SI23: Severity of patch artifacts across different overlapping pixel sizes, visualized using the mean absolute Sobel-X gradient of surface temperature field as a function of horizontal pixel position.
Ensemble of experts and over-fitting
As discussed in Section A.3.2 the coarse-resolution cBottle-3d is prone to over-fitting the largest modes of the data and propose a regularized ensemble of experts approach as a fix. This was initially pointed out to us by astute readers of an early version of this manuscript, who noticed the over-fitting in the Northern Annular Mode (see 4b). Intriguingly, this also improves the sample translation abilities demonstrated in Fig 5. The translated sample by the regularized approach Fig SI24c much more closely resembles the original sample ( Fig. SI24a). On the other hand, the over-fit model SI25 hallucinates an entirely new storm system in the south pacific and elsewhere.
a. ERA5 b. ICON (overfit) c. ICON (regularized)
Refer to caption Refer to caption Refer to caption
c. d. e.
Refer to caption Refer to caption Refer to caption
Figure SI24: Impact of ensemble of experts regularization on the data-set translation example shown in Fig. 5. (a-c) dBR. (a) The original ERA5 dBR field. This sample translated to ICON by (b) the over-fit denoiser trained to 9.8M images and (c) the regularized ensemble of experts approach copied from Fig 5d. (d-e) is the same but for column integrated liquid water
Figure SI25: Diffusion loss versus $\sigma$ for different training iterations for the training and a subset of the validation set. From upper left, $\sigma\in(0,0.1]$, (0.1, 1], (10, 100], (100, 1000]. The x-axis is the training progress in 1000s of samples.